OpenStudy (anonymous):

The square root of 16 to the fourth power. The square root of 16 to the fourth power. @Mathematics

OpenStudy (saifoo.khan):

256

OpenStudy (anonymous):

wait so is it 16^4 or 16^2

OpenStudy (saifoo.khan):

\[\Large \sqrt{16^4} = \sqrt{2^{16}}= 2^8=256\]

OpenStudy (anonymous):

ok